OpenStudy (anonymous):

If y varies directly as x,and y-15 as x=-3 find y for the x - value of 4

OpenStudy (cggurumanjunath):

try similarly !

OpenStudy (cggurumanjunath):

y=kx ( y varies directly as x )

OpenStudy (cggurumanjunath):

when y=15 and x=-3 find k

OpenStudy (cggurumanjunath):

is y +15 or -15 ? @pjpkap

OpenStudy (anonymous):

-15

OpenStudy (cggurumanjunath):

ok

OpenStudy (cggurumanjunath):

when y=-15 and x=-3 find k

OpenStudy (cggurumanjunath):

@pjpkap

OpenStudy (texaschic101):

I have a question....if it varies directly, can it be solved with proportions ?

OpenStudy (cggurumanjunath):

y=kx -15=k*-3

OpenStudy (cggurumanjunath):

now can u find k @pjpkap

OpenStudy (cggurumanjunath):

how @texaschic101

OpenStudy (cggurumanjunath):

can u show ur steps @texaschic101 ?

OpenStudy (anonymous):

I just dont get it very well?

OpenStudy (texaschic101):

-15/-3 = 4/x can it be solved that way ??

OpenStudy (cggurumanjunath):

y=kx -15=k*-3 now can u find k @pjpkap

OpenStudy (anonymous):

no

OpenStudy (cggurumanjunath):

-15=k*-3

OpenStudy (cggurumanjunath):

-15/-3 = y/4 @texaschic101

OpenStudy (texaschic101):

yes...thats it...I wrote it wrong

OpenStudy (texaschic101):

can you solve it that way ?

OpenStudy (cggurumanjunath):

yes u can ! @texaschic101

OpenStudy (texaschic101):

if it is indirectly would you use indirect proportion ?

OpenStudy (cggurumanjunath):

but sometimes variation problems involves a lot of variables ,then ur method won't work.

OpenStudy (cggurumanjunath):

-15=k*-3 @pjpkap

OpenStudy (texaschic101):

oh..I see...I have trouble with these, so I was just asking. Thanks for explaining :)

OpenStudy (cggurumanjunath):

divide both sides by -3 .@pjpkap

OpenStudy (anonymous):

Is the following a fair restatement of the problem? y varies directly as x. When y = 15, then x = -3. Find y when x = 4.

OpenStudy (texaschic101):

I thought y was -15

OpenStudy (cggurumanjunath):

yes y=-15 only as told by @pjpkap

OpenStudy (anonymous):

It is -15

OpenStudy (cggurumanjunath):

-15=k*-3 @pjpkap divide both sides by -3 .@pjpkap

OpenStudy (cggurumanjunath):

(-15/-3)=(k*-3)/(-3)

OpenStudy (cggurumanjunath):

@pjpkap

OpenStudy (cggurumanjunath):

now can u tell me the value of k ?

OpenStudy (anonymous):

Im here just confused

OpenStudy (cggurumanjunath):

-15=k*-3 @pjpkap divide both sides by -3 .@pjpkap

OpenStudy (texaschic101):

can you solve it if it is written like this : -15 = -3k ...pjpkap ?

OpenStudy (anonymous):

k=-15/3 k=-5 ??

OpenStudy (texaschic101):

-15/-3 = what

OpenStudy (cggurumanjunath):

k=-15/-3 ; k=5

OpenStudy (cggurumanjunath):

isn't it @pjpkap ?

OpenStudy (anonymous):

yes

OpenStudy (cggurumanjunath):

ok

OpenStudy (cggurumanjunath):

y=kx k=5 x=4

OpenStudy (cggurumanjunath):

now can u find y by substituting in y=kx

OpenStudy (cggurumanjunath):

@pjpkap

OpenStudy (texaschic101):

nice explaining CGGURUMANJUNATH :)

OpenStudy (cggurumanjunath):

:)

OpenStudy (cggurumanjunath):

thank u . and u too. @texaschic101

OpenStudy (cggurumanjunath):

y=kx k=5 x=4 now can u find y by substituting in y=kx @pjpkap

OpenStudy (anonymous):

Im trying ,I have dyslexia and all my numbers are getting jumbled.

OpenStudy (cggurumanjunath):

y=kx k=5 x=4 y=kx y=5*4

OpenStudy (cggurumanjunath):

how much is 5 times 4 @pjpkap

OpenStudy (anonymous):

20

OpenStudy (cggurumanjunath):

u got it .

OpenStudy (cggurumanjunath):

:)

OpenStudy (cggurumanjunath):

y=20

OpenStudy (cggurumanjunath):

that is your required value.

OpenStudy (anonymous):

thank you very much, i wish my teacher had as much patience with me.

OpenStudy (cggurumanjunath):

y=20 is the final answer .

OpenStudy (cggurumanjunath):

:)